OpenStudy (anonymous):

Solve the equation sin (x° - 20°) = cos 42° for x, where 0 < x < 90 A. 62 B. 22 C. 68 D. 28

OpenStudy (therealmeeeee):

ok cos(42)=sin(90-42)=sin(48) sin(x-20)=sin(48) x-20=48 x=48+20 x=68 degrees

OpenStudy (therealmeeeee):

this make sense

OpenStudy (anonymous):

yea kind of... thanks though

OpenStudy (therealmeeeee):

can u click best response

OpenStudy (therealmeeeee):

ty

OpenStudy (anonymous):

lol np